40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9068     Entry time: Tue Aug 27 02:18:28 2013
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRMI, DRMI sensing matrices 

I have made a measurement of the PRMI and the DRMI sensing matrices. 

Keiko pointed out to me in an email a little while ago that I wasn't zeroing elements in the oscillator drive matrix after using them, so I was effectively driving all the degrees of freedom at once, which is why some of the recent sensing matrices looked a little bullshitty.  Anyhow, I put in a few lines to zero that row of the LSC  output matrix, so that we don't do that any more. 

PRMI sensing matrix:

SensMatMeas_26Aug2013_withErrorBars.png

DRMI sensing matrix, first-ever measurement, after the optic flipping / recent locking success:

SensMatMeas_26Aug2013_DRMI_firstEverMeasurement.png

Note that we don't have any error bars in the DRMI case, since the IFO fell out of lock during the error bar measurements.  So, we got the "real" data from the degrees of freedom, but not extra data for making error bars.  Also, the MICH / SRCL coupling hasn't been balanced out in the output matrix yet, but since the notches are engaged in the degrees of freedom during this measurement, that shouldn't be a significant effect.

To get the DRMI sensing matrix measured, I added the SRM's actuator calibration to SensMatDefinitions.py (data from elog 5637).  I also created a new file runDRMI_sens, to be the equivalent of runPRMI_sens.  In the new runDRMI_sens, I reduced the actuation from the oscillator by a factor of 10.  I had several attempts at higher oscillator amplitudes that kept kicking the IFO out of lock.

The DRMI was pretty good, but I wasn't getting ~10s of minutes like Koji was on Friday.  I also wasn't able to engage all of the FM triggers that he was.  The 10-30 Hz seismic BLRMS is a little higher than a usual night, but other than that, seismic looks pretty quiet.

My settings for the night:

LSC input matrix:  +0.1*REFL55Q = MICH, -0.125*REFL11I = PRCL, +1.00*REFL55I = SRCL.

Filter settings:  MICH, PRCL, SRCL all had FM4,5 always on.  MICH had FM2,3 triggered.  PRCL had FM2,3,6 triggered.  SRCL had FM2 triggered.  In particular, engaging FM 6 for MICH or SRCL made some loud low-ish frequency oscillation.  Engaging anything other than FM2 for SRCL kicked the IFO out of lock. 

Gains:  MICH = -0.800, PRCL = +0.050, SRCL = -0.100

Triggering:  All triggered on POP22I, upper = 50, lower = 10 (lower = 25 for SRCL).

FM trigger thresholds: MICH on = 35, off = 2, delay = 2 sec.  PRCL on = 35, off = 2, delay = 0.5 seconds.  SRCL on = 80, off = 25, delay = 5 sec.

Power normalization:  None, for any degree of freedom.

LSC Output matrix:  MICH = -0.267 for PRM, +0.50 for BS.  PRCL = +1.0 for PRM.  SRCL = +1.0 for SRM. 

LSC SUS filters:  BS, PRM, SRM all had FM1,2,3,6 engaged for the BS, PRM and SRM violin filters, as well as the 3rd order harmonic for one of them.

Other notes: 

I tried locking the SRMI, so that I could do the same kind of actuator calibration that Koji did for the PRMI in elog 8816, but was unsuccessful.  I checked optickle, and found that for REFL 55 I&Q locking, MICH and SRCL keep the same signs for SRMI as DRMI.  Also, for both, the optical response is a factor of ~15 lower for SRMI than DRMI, so the gains should be higher by a factor of 15 for both MICH and SRCL.  I think my big problem here is that I don't have anything to trigger on.  There isn't any signal to speak of in the POP PDs, with the PRM misaligned.  Hopefully we'll have AS110 shortly, and that will help.

I updated the IFO Configure restore scripts to our latest versions of locking.  I have also tested them, and restoring the Michelson, PRMI and DRMI all seem to work. (MICH restores to locking with AS55Q.  PRMI restores to locking with REFL165 I&Q.  DRMI restores to the settings noted above in this entry.)  The X and Y arm restores have been working, and I have been using them (semi-)regularly since I announced them in elog 8433 back in April.  Still to-do though:  Add PRCL ASC to the PRMI up script, and make the dither options work for at least the arms and PRM.  (Just need to point the drop down menu options to the new ASS scripts.)

 

 

ELOG V3.1.3-